# Information conservation in de Sitter tunneling

### v1

@@ 5 Conclusion and Discussion
In the above calculation, we discussed the three most general cases of asymptotic de Sitter spacetime. The charged and magnetic particles tunnel into the magnetically charged Reissner-Nordström de Sitter black hole (the most general case of a static black hole), the Kerr-Newman-Kasuya de Sitter black hole (the most general case of a rotating black hole), and Bardeen de Sitter black hole (black hole without singularities). We calculated the radiation spectra of particles exiting event horizon of a black hole and entering cosmic horizon using the Parikh-Wilczek method. Our results indicate that for these three most general cases, the particle radiation spectrum deviates from the pure thermal spectrum, in accordance with the unitary principle and supporting information conservation. That is, information conservation is generally valid for general asymptotic de Sitter space-times. The fundamental reason for information conservation in the Parikh-Wilczek framework is the requirement that radiation process of particles is a reversible quasi-static process. For reversible processes, entropy is conserved, and information is naturally conserved as well.
Acknowledgements.
The authors would like to thank Prof. Jian Jing and his student LiuBiao Ma and Zheng Wang from the Department of Physics , Beijing University of Chemical Technology for their valuable comments and suggestions during the completion of this manuscript.

### v2

@@ 7 Conclusion and Discussion
In the above calculation, we discussed the three most general cases of asymptotic de Sitter spacetime. The charged and magnetic particles tunnel into the magnetically charged Reissner-Nordström de Sitter black hole (the most general case of a static black hole), the Kerr-Newman-Kasuya de Sitter black hole (the most general case of a rotating black hole), and Bardeen de Sitter black hole (black hole without singularities). We calculated the radiation spectra of particles exiting event horizon of a black hole and entering cosmic horizon using the Parikh-Wilczek method. Our results indicate that for these three most general cases, the particle radiation spectrum deviates from the pure thermal spectrum, in accordance with the unitary principle and supporting information conservation. That is, information conservation is generally valid for general asymptotic de Sitter space-times. The fundamental reason for information conservation in the Parikh-Wilczek framework is the requirement that radiation process of particles is a reversible quasi-static process. For reversible processes, entropy is conserved, and information is naturally conserved as well. Our discoveries not only strengthen the coherence of quantum mechanics within cosmological horizons but also offer essential contributions to the burgeoning dS/CFT correspondence.
[…]
Acknowledgements The authors are grateful to anonymous reviewers for their very crucial and detailed comments. The authors would like to thank Prof. Jian Jing and his student LiuBiao Ma and Zheng Wang from the Department of Physics , Beijing University of Chemical Technology for their valuable comments and suggestions during the completion of this manuscript.